### 10 Mar 2026 v1

# X-GS: An Extensible Open Framework Unifying 3DGS Architectures with Downstream Multimodal Models

Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful technique for novel view synthesis, subsequently extending into numerous spatial AI applications. However, most existing 3DGS methods are isolated, focusing on specific domains such as online SLAM, semantic enrichment, or 3DGS for unposed images. In this paper, we introduce X-GS, an extensible open framework that unifies a broad range of techniques to enable real-time 3DGS-based online SLAM enriched with semantics, bridging the gap to downstream multimodal models. At the core of X-GS is a highly efficient pipeline called X-GS-Perceiver, capable of taking unposed RGB (or optionally RGB-D) video streams as input to co-optimize geometry and poses, and distill high-dimensional semantic features from vision foundation models into the 3D Gaussians. We achieve real-time performance through a novel online Vector Quantization (VQ) module, a GPU-accelerated grid-sampling scheme, and a highly parallelized pipeline design. The semantic 3D Gaussians can then be utilized by vision-language models within the X-GS-Thinker component, enabling downstream tasks such as object detection, zero-shot caption generation, and potentially embodied tasks. Experimental results on real-world datasets showcase the efficacy, efficiency, and newly unlocked multimodal capabilities of the X-GS framework.
X-GS: An Extensible Open Framework Unifying 3DGS Architectures with Downstream Multimodal Models
Yueen Ma1, Irwin King1 Department of Computer Science and Engineering, The Chinese University of Hong Kong1 {yema21, king}@cse.cuhk.edu.hk
1 Introduction
Figure 1: X-GS is an extensible open framework that unifies previously isolated domains, including pose-free 3DGS, 3DGS-based SLAM, Semantic GS, and VLMs for 3DGS. X-GS-Perceiver achieves real-time 3DGS-based online SLAM with semantics using an online VQ module, GPU-accelerated grid-sampling, and highly parallelized scheduling. X-GS-Thinker bridges the resulting semantic 3DGS representation with downstream multimodal models to execute complex, language-driven tasks such as open-vocabulary 3D object detection, caption generation, and potentially embodied tasks.
3D Gaussian Splatting (3DGS) Kerbl et al. (2023) has achieved groundbreaking success in real-time novel view synthesis, catalyzing rapid advancements across several distinct research directions. To eliminate the reliance on computationally expensive offline Structure-from-Motion (SfM) pipelines like COLMAP, recent methods Fu et al. (2024) have successfully reconstructed scenes directly from unposed image sequences by relaxing camera pose requirements. Furthermore, the advent of 3DGS-based SLAM systems Matsuki et al. (2024) has improved upon traditional point-cloud-based methods, allowing the mapped scene to retain significantly richer visual information. Extending beyond pure appearance, Semantic-GS Li et al. (2025c) endows the 3D Gaussians with explicit semantic feature channels alongside their standard parameters to enable complex scene understanding. Finally, bridging the gap between spatial representations and multimodal reasoning, the latest works have developed Vision-Language Models (VLMs) Thai et al. (2025) that natively utilize 3D Gaussians as direct inputs rather than relying solely on 2D images.
Despite this rapid progress, these advancements typically operate in isolation. To combine their respective strengths, we propose the first work to unify these disparate 3DGS techniques from a comprehensive framework perspective: the X-GS framework. Standing for an eXtensible open framework for 3DGS, the “X” serves as a placeholder prefix to denote distinct 3DGS paradigms, such as Pose-Free-GS, SLAM-GS, Semantic-GS, and VLM-for-GS. This suggests that future developments in these different directions can be assimilated by X-GS for continued improvement, directly showcasing its extensibility. Built upon this unified foundation, X-GS achieves real-time 3DGS-based online SLAM with rich semantic enrichment while simultaneously bridging the gap to downstream VLMs.
Table 1: Comparison of X-GS with representative 3DGS methods.
Method RGB-only RGB-D Pose-free Online-SLAM Real-time Semantics VLM-Integration
Offline 3DGS & Semantic Fields
[…]
Online 3DGS SLAM
[…]
VLMs for 3DGS
[…]
X-GS (Ours) $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Although X-GS is built upon existing 3DGS algorithms, integrating them for real-time performance is a non-trivial process. To achieve this, we introduce X-GS-Perceiver, our dedicated semantic mapping module, which is driven by three key efficiency techniques: an online vector quantization (VQ) module, a grid-sampling scheme, and a meticulously designed pipeline that fully leverages parallelism. The VQ module drastically reduces the dimensionality of the semantic channels stored in each 3D Gaussian. While previous works have utilized VQ modules in 3DGS, they were limited to offline processing. We overcome this by introducing a VQ module with Exponential Moving Average (EMA) updates to support continuous online learning. Additionally, because 3D Gaussians usually project to areas rather than single pixels on the rendered image, we introduce a grid-sampled semantic supervision scheme. This is paired with a custom GPU kernel that executes only the minimal calculations required for the subsampled pixels. This technique maintains semantic map quality while achieving significant speedup and memory load reduction. At a systemic level, X-GS-Perceiver schedules these individual components using highly parallelized strategies, such as early VQ codebook updates and grid-sampled target prefetching.
Because X-GS can accommodate various vision foundation models, it readily supports multiple downstream multimodal tasks through the X-GS-Thinker component. By combining SAM Kirillov et al. (2023) and CLIP Radford et al. (2021), X-GS-Perceiver receives object-level semantics. The resulting 3D Gaussians thus encode these semantics, enabling X-GS-Thinker, which in this case is based on the CLIP text encoder, to perform text-prompted 3D object detection in the mapped scene. If we integrate the vision tower of a VLM Liu et al. (2023), the 3DGS semantics can also be fed into the model’s backbone, acting as the X-GS-Thinker, to generate scene captions. Potentially, by connecting these semantic features to a vision-language-action (VLA) model Brohan et al. (2023), its action generation module can serve as the X-GS-Thinker to tackle embodied tasks. This suggests the additional extensibility of the X-GS framework to downstream multimodal models.
In summary, the main contributions are:
  • We propose X-GS, an extensible open framework that unifies previously isolated domains—such as 3DGS-based online SLAM, pose-free 3DGS, semantic 3DGS, and 3DGS-based VLMs—into a cohesive system.
  • We introduce X-GS-Perceiver, equipped with three core optimization techniques: an online VQ module with EMA updates, a GPU-accelerated grid-sampling scheme, and a highly parallelized scheduling design, collectively ensuring real-time performance.
  • We demonstrate its additional extensibility by introducing the X-GS-Thinker component, which bridges the resulting semantic 3DGS with downstream multimodal models to enable tasks such as 3D object detection, scene captioning, and embodied tasks.
2 Related Work
Figure 2: Overview of the X-GS framework. X-GS-Perceiver synergizes a memory-efficient Vector Quantization (VQ) module, grid-based semantic supervision, and an asynchronous parallelized pipeline to perform SLAM and distill semantics simultaneously in an online fashion, operating in real time at $\sim$15 FPS. As an open framework, it accommodates both RGB-only and RGB-D inputs, and can flexibly integrate various Vision Foundation Models (VFMs). Furthermore, the X-GS-Thinker component is extensible to different multimodal models, enabling a wide range of downstream tasks.
2.1 3DGS-based Online SLAM
Standard 3D Gaussian Splatting (3DGS) algorithms Kerbl et al. (2023) rely on offline multi-view image sequences and COLMAP initialization. While subsequent methods like CF-3DGS Fu et al. (2024) eliminate the need for pre-computed poses, they still require offline processing. Recently, approaches such as MonoGS Matsuki et al. (2024), GS-SLAM Yan et al. (2024), and Gaussian-SLAM Yugay et al. (2023) have achieved online, real-time SLAM using 3DGS. However, these systems focus exclusively on 3D reconstruction and camera tracking, ignoring high-level semantic scene understanding. To enrich 3DGS representations, Feature 3DGS Zhou et al. (2024) and LangSplat Qin et al. (2024); Li et al. (2025c) distill dense semantic features from vision foundation models (e.g., SAM or CLIP) directly into the 3D Gaussian field, enabling language-driven queries and zero-shot 3D object detection. Nevertheless, these semantic methods rely heavily on precise, pre-computed camera poses and are strictly designed for offline mapping, inherently limiting their applicability in dynamic or autonomous environments.
2.2 Semantic 3DGS-based SLAM
Existing semantic 3DGS SLAM systems, such as LEGO-SLAM Lee et al. (2025) and OpenMonoGS-SLAM Yoo et al. (2025), attempt online semantic mapping but suffer from rigid architectural bottlenecks. LEGO-SLAM strictly requires RGB-D input, whereas OpenMonoGS-SLAM operates solely on monocular video—lacking the flexibility to utilize depth maps even when available. Furthermore, OpenMonoGS-SLAM fails to achieve real-time performance. X-GS natively overcomes all of these limitations.
2.3 VLMs for 3DGS
Recent literature Li et al. (2024) demonstrates the immense value of explicitly integrating 3D representations with Vision-Language Models (VLMs) for downstream reasoning. For instance, UniGS Li et al. (2025b) aligns optimized 3D Gaussians with textual spaces for multimodal contrastive learning, while SplatTalk Thai et al. (2025) and ChatSplat Chen et al. (2024) build VLMs that take 3DGS as visual input. However, current VLMs for 3DGS are restricted entirely to static, offline scenes. X-GS unifies these disjointed paradigms into a single online framework.
Table 1 provides a comparison of key capabilities among representative 3DGS methods. To the best of our knowledge, X-GS is the first framework to fulfill all the listed requirements online.
3 Method
We propose X-GS, an extensible open framework designed to unify previously isolated 3DGS advancements. The architecture is divided into two primary subsystems. First, the perception module, X-GS-Perceiver, ingests unposed monocular or RGB-D video streams to continuously co-optimize the 3D Gaussian map and camera poses, while distilling high-dimensional foundation model features into the Gaussians as semantic channels. To ensure real-time performance, it utilizes online vector quantization, grid-sampling, and parallel scheduling. Second, different multimodal models can serve as the X-GS-Thinker. By utilizing the enriched 3DGS representation produced by X-GS-Perceiver, the Thinker unlocks advanced capabilities such as text-prompted 3D object detection, zero-shot scene captioning, and embodied tasks.
3.1 Preliminaries
3D Gaussian Splatting and Feature Rendering.
3DGS represents scene radiance (geometry and appearance) using a collection of 3D Gaussians. Each Gaussian is parameterized by $\Theta=\{\mu,\Sigma,\alpha,c\}$, representing its 3D center $\mu$, covariance $\Sigma$, base opacity $\alpha$, and color $c$. To compute the expected color $C(x)$ for a pixel $x$ from a specific camera pose $\mathbf{T}\in\text{SE}(3)$, a set of $N$ overlapping 3D Gaussians are transformed by $\mathbf{T}$ into camera space, projected into 2D screen coordinates, sorted by depth, and composited via front-to-back $\alpha$-blending:
$C(x\mid\mathbf{T},\Theta)=\sum_{i=1}^{N}c_{i}\alpha^{\prime}_{i}\prod_{j=1}^{i-1}(1-\alpha^{\prime}_{j})$ (1)
where $c_{i}$ is the color and $\alpha^{\prime}_{i}$ is the evaluated 2D footprint opacity of the $i$-th Gaussian. Crucially, $\alpha^{\prime}_{i}$ and the depth-sorting order are explicit functions of the camera pose $\mathbf{T}$ via the view transformation and perspective projection, rendering the entire rasterization process differentiable with respect to both $\mathbf{T}$ and $\Theta$.
Beyond RGB appearance, this rasterization process natively extends to high-dimensional semantic features. By substituting the 3-dimensional color $c_{i}$ with a $D$-dimensional feature vector, the exact same $\alpha$-blending formulation in (1) renders a dense 2D semantic feature map $F(x\mid\mathbf{T},\Theta)$. This computationally pairs the discrete 3D semantic states of the Gaussians with their corresponding 2D spatial observations.
3DGS-based SLAM.
The online SLAM pipeline Matsuki et al. (2024) operates via two continuous, concurrent threads: tracking and mapping. During tracking, camera poses are estimated through direct optimization against the explicit 3D Gaussian representation. The pose $\mathbf{T}_{t}$ for an incoming frame $t$ is initialized using a constant velocity kinematic model. Leveraging the differentiability of (1), $\mathbf{T}_{t}$ is then iteratively refined via gradient descent by holding the map parameters $\Theta$ fixed and minimizing the tracking loss $\mathcal{L}_{\text{track}}$ against the captured ground-truth color image $I_{t}$ and optional depth map $D_{t}$:
$\mathbf{T}_{t}^{*}=\underset{\mathbf{T}_{t}}{\arg\min}\ \mathcal{L}_{\text{track}}(I_{t},D_{t}\mid\mathbf{T}_{t},\Theta)$ (2)
Concurrently, the mapping thread refines the global scene representation by holding the camera poses fixed and optimizing the Gaussian parameters $\Theta$. This optimization is performed over a sliding window of keyframes $\mathcal{K}$, which is dynamically maintained by inserting new frames that provide sufficient novel viewpoints, while evicting redundant older views to preserve a maximum capacity of $N_{win}$. The mapping thread refines $\Theta$ by minimizing the mapping loss $\mathcal{L}_{\text{map}}$ across this active window of ground-truth images $I_{k}$ and depth maps $D_{k}$:
$\Theta^{*}=\underset{\Theta}{\arg\min}\sum_{k\in\mathcal{K}}\mathcal{L}_{\text{map}}(I_{k},D_{k}\mid\Theta,\mathbf{T}_{k})$ (3)
Dynamic mapping heuristics, such as geometry-guided Gaussian insertion and opacity-based pruning, are utilized alongside these objectives to maintain a structurally clean topological map.
3.2 X-GS-Perceiver
We develop a set of modules that add semantic enrichment capabilities to the online geometric and photometric optimization pipeline. Naively attaching semantic features to 3D Gaussians has proven to be heavily memory- and compute-intensive. To ensure our pipeline still fulfills the real-time requirement, we devise three key speedup techniques: an online vector quantization (VQ) module, a GPU-accelerated grid-sampling scheme, and a meticulously designed pipeline that fully leverages highly parallelized scheduling.
Online Vector Quantization Module.
We represent the semantic state of each Gaussian by logits over a shared codebook, inspired by Li et al. (2025c). Let $N_{g}$ denote the number of Gaussians, $K$ the number of codewords, and $D$ the feature dimension. The shared codebook parameters are
$\mathbf{E}\in\mathbb{R}^{K\times D}.$ (4)
Each Gaussian $i$ stores a learnable logit vector
$\mathbf{z}_{i}\in\mathbb{R}^{K}.$ (5)
We convert these logits into mixture weights using a softmax over all codebook entries:
$\mathbf{w}_{i}$ $=\operatorname{softmax}(\mathbf{z}_{i}),$ (6)
$w_{ik}$ $=\frac{\exp(z_{ik})}{\sum_{k^{\prime}=1}^{K}\exp(z_{ik^{\prime}})}.$
The decoded semantic feature of Gaussian $i$ is then obtained by a weighted sum over the shared codebook:
$\hat{\mathbf{f}}_{i}$ $=\mathbf{E}^{\top}\mathbf{w}_{i}$ (7)
$=\sum_{k=1}^{K}w_{ik}\,\mathbf{e}_{k}\in\mathbb{R}^{D}.$
This parameterization separates Gaussian-specific semantic coefficients from the shared semantic atoms in the codebook, yielding a compact representation that can be optimized jointly with the Gaussian scene.
We update the shared codebook online using vector quantization with Exponential Moving Averages (EMA). Given a set of observed semantic features $\{\mathbf{x}_{n}\}_{n=1}^{N}$, each feature is assigned to its nearest codeword:
$a_{n}=\arg\min_{k\in\{1,\dots,K\}}\left\|\mathbf{x}_{n}-\mathbf{e}_{k}\right\|_{2}.$ (8)
We then accumulate the assignment counts and feature sums,
$c_{k}$ $=\sum_{n=1}^{N}\mathbb{I}\!\left[a_{n}=k\right],$ (9)
$\mathbf{s}_{k}$ $=\sum_{n=1}^{N}\mathbb{I}\!\left[a_{n}=k\right]\mathbf{x}_{n},$
and update the EMA buffers
$N_{k}$ $\leftarrow\lambda N_{k}+(1-\lambda)c_{k},$ (10)
$\mathbf{M}_{k}$ $\leftarrow\lambda\mathbf{M}_{k}+(1-\lambda)\mathbf{s}_{k},$
where $\lambda\in[0,1)$ is the EMA decay. The updated codeword is
$\mathbf{e}_{k}\leftarrow\frac{\mathbf{M}_{k}}{N_{k}+\varepsilon},$ (11)
with a small constant $\varepsilon$ for numerical stability. In this way, the per-Gaussian logits are learned by gradient descent, while the shared codebook tracks the evolving feature distribution through EMA continuous online learning.
Grid-Sampled Semantic Supervision.
Applying dense semantic supervision at every image pixel is unnecessarily expensive during mapping, because 3D Gaussians project to areas rather than single pixels on the rendered image. We therefore supervise the rendered semantic features on a regular stride-offset grid in the image plane. Let the full image domain be
$\Omega=\{0,\dots,H-1\}\times\{0,\dots,W-1\},$ (12)
and let the grid stride be $s\geq 1$ with offset $\mathbf{o}=(o_{h},o_{w})$, where $0\leq o_{h},o_{w}<s$. The sampled grid is
$\Omega_{s,\mathbf{o}}=\left\{(u,v)\in\Omega\;\middle|\;\right$. $u\equiv o_{h}\!\!\!\pmod{s}$, (13)
$\left.v\equiv o_{w}\!\!\!\pmod{s}\right\}$.
Equivalently, each sampled location can be written as
$(u,v)=(o_{h}+ms,\;o_{w}+ns),$ (14)
for integer indices $m,n$. The resulting compact feature-map resolution is
$H_{s}$ $=\max\!\left(0,\;1+\left\lfloor\frac{H-1-o_{h}}{s}\right\rfloor\right),$ (15)
$W_{s}$ $=\max\!\left(0,\;1+\left\lfloor\frac{W-1-o_{w}}{s}\right\rfloor\right).$
Let $\hat{\mathbf{F}}\in\mathbb{R}^{D\times H\times W}$ denote the dense rendered semantic feature field. Conceptually, grid sampling restricts this field to the grid $\Omega_{s,\mathbf{o}}$, yielding the compact prediction
$\hat{\mathbf{G}}_{:,m,n}=\hat{\mathbf{F}}_{:,\,o_{h}+ms,\,o_{w}+ns}\in\mathbb{R}^{D}.$ (16)
To maximize efficiency, we pair this scheme with a custom GPU kernel optimized to execute only the minimal pixel-level calculations required for the subsampled grid locations. The rasterizer directly outputs this compact prediction, entirely avoiding the computational and memory burdens of forming a dense semantic feature map before subsampling.
We construct the target feature map $\mathbf{G}^{\star}\in\mathbb{R}^{D\times H_{s}\times W_{s}}$ on the same grid, together with a binary validity mask
$\mathbf{V}\in\{0,1\}^{1\times H_{s}\times W_{s}}.$ (17)
The semantic supervision is then applied only on valid sampled locations:
$\mathcal{L}_{\mathrm{grid}}=$ $\lambda_{\mathrm{sem}}\left[\mathcal{L}_{\cos}\!\left(\mathbf{V}\odot\hat{\mathbf{G}},\;\mathbf{V}\odot\mathbf{G}^{\star}\right)\right$. (18)
$\quad\left.+\,\mathcal{L}_{1}\!\left(\mathbf{V}\odot\hat{\mathbf{G}},\;\mathbf{V}\odot\mathbf{G}^{\star}\right)\right],$
where $\odot$ denotes element-wise masking. The validity mask $\mathbf{V}$ serves two vital purposes. Primarily, it acts as a filter for unannotated or background regions that yield meaningless zero-vectors in the target grid $\mathbf{G}^{\star}$, preventing the 3D Gaussians from incorrectly learning empty features. Secondarily, it masks out padded, out-of-bounds coordinates that arise from maintaining rigid batched tensor shapes at large grid offsets, ensuring the loss is strictly evaluated on robust semantic data.
When $s=1$, the formulation reduces to standard dense semantic supervision. However, for a generic stride $s>1$, the custom kernel correctly reduces the number of supervised pixels by a factor of $s^{2}$. This explicit reduction yields a proportional $s^{2}\times$ savings in both the memory bandwidth required for rasterizing the deep semantic features and the computational overhead of the loss evaluation, keeping the entire optimization comfortably within real-time constraints.
Parallel Pipeline Architecture.
To integrate these individual components into a highly efficient workflow, we employ a heavily parallelized system architecture, as illustrated in Figure 2. The VQ codebook update is executed as soon as the Vision Foundation Model (VFM) completes encoding an incoming keyframe. Simultaneously, we initiate a “Grid-Sampled Target Prefetching” operation on the full-resolution semantic map generated by the VFM. Because operations across different spatial grid offsets are mutually independent, we allocate multiple background workers to extract and buffer the grid-sampled targets well in advance of the actual optimization loop.
Crucially, to maintain system stability and computational efficiency, we strictly decouple the geometry and appearance updates from the semantic updates. During the semantic optimization phase, all foundational parameters of the 3D Gaussians (i.e., position, scale, rotation, opacity, and color) remain frozen. Conversely, the learned semantic logits and VQ codebooks are strictly excluded from the base tracking and mapping steps. By synergizing the memory-efficient VQ module, the reduced computational footprint of our grid-based semantic supervision, and this multi-threaded alternating optimization schedule, our system achieves comprehensive semantic enrichment while consistently maintaining real-time performance at $\sim$15 FPS.
Figure 3: Qualitative results of X-GS on scene reconstruction and semantic distillation. From left to right: Ground Truth (GT) RGB, Rendered RGB, GT Semantic Map (from VFMs, SAM + CLIP), Rendered Semantic Map, and an open-vocabulary Object Detection example.
3.3 X-GS-Thinker
While the X-GS-Perceiver handles the efficient construction of 3D semantic Gaussians, the X-GS-Thinker operates atop this framework to interpret and leverage these learned representations for downstream multimodal tasks. Because the semantic fields mapped onto the 3D Gaussians are distilled directly from prominent vision foundation models (e.g., CLIP, SAM, or SigLIP), the X-GS-Thinker remains highly extensible, adapting seamlessly to various multimodal network architectures.
Contrastive VLM.
For text-prompted 3D object detection, the X-GS-Thinker relies on a contrastive vision-language model, such as OpenCLIP. Crucially, rather than rendering 2D semantic feature maps, we localize user-specified concepts by querying the 3D scene representation directly. First, the text query and a standard set of generic negative phrases (e.g., “object”, “things”, “stuff”, “texture”) are encoded and $L_{2}$-normalized to construct a contrastive subspace. Then, for each 3D Gaussian in the scene, we decode its continuous semantic embedding by multiplying its learned mixture weights with the shared codebook. We compute a prompt-conditioned relevance score natively in 3D by contrasting each Gaussian’s decoded vector against the encoded text phrases, applying a temperature-scaled softmax to penalize features that strongly correlate with negative structural noise. By thresholding these explicit 3D relevancy queries, we generate a prompt-specific Gaussian mask. The surviving subset of Gaussians can then be rasterized from any arbitrary viewpoint to isolate the target object, effectively performing deterministic, view-independent open-vocabulary 3D segmentation without requiring explicit bounding box inferences.
Generative VLM.
To achieve higher-level scene understanding, the decoded 3D Gaussian semantic features can be routed directly into the vision tower of a Generative Vision-Language Model (VLM). However, passing the full set of $N_{g}$ Gaussians into a generative model is computationally prohibitive and yields highly redundant tokens. Thus, we condense the scene into a compact sequence of $M$ informative tokens using an Entropy-Adaptive Gaussian Sampling strategy.
Rather than relying on naive spatial down-sampling, this strategy leverages the information-theoretic uncertainty inherent in each Gaussian’s learned semantic state. Specifically, we evaluate the Shannon entropy of the semantic assignment probabilities produced by the VQ mixture logits. Gaussians exhibiting high entropy—often corresponding to ambiguous object boundaries, interactive structural frontiers, or geometrically dense regions—are actively prioritized. In contrast, Gaussians with low entropy tightly map to a single categorical cluster, indicating redundant, homogeneous backgrounds (like large flat floors or empty walls). We run a continuous Top-$M$ sorting algorithm over these calculated entropy scores, explicitly extracting only the subset of Gaussians exhibiting peak semantic ambiguity. This deterministic sampling naturally distills out redundant backgrounds while preserving critical semantic boundaries and object structures. The resulting token sequence seamlessly bridges the continuous geometric scene representation with the VLM’s discrete context window for complex reasoning tasks, such as 3D visual question answering (VQA) and zero-shot scene captioning.
Potential Embodied AI Application.
Finally, as a future direction for embodied AI applications, the X-GS-Thinker can be configured to interface with a Vision-Language-Action (VLA) model Brohan et al. (2023). By feeding language-aligned 3D geometric features directly into a VLA, our system provides the real-time spatial information required to support embodied tasks.
4 Experiments
4.1 Implementation Details
Our 3DGS-based online SLAM pipeline for geometry and appearance builds upon MonoGS Matsuki et al. (2024), which executes online SLAM via two concurrent threads: a tracking thread for camera pose estimation and a mapping thread for optimizing the geometry and appearance of the 3D Gaussians. During the online codebook learning stage Li et al. (2025c), the EMA VQ codebook is instantiated with $K=256$ independent semantic codes. The EMA update incorporates a decay momentum of $\lambda=0.96$ and a temperature of $\tau=0.7$ to smoothly adapt the cluster centers to the sequential frame stream.
To extract object boundaries, we employ SAM Kirillov et al. (2023), utilizing its base backbone to obtain semantic region masks. To inject open-vocabulary capabilities into the 3D Gaussians, we utilize CLIP (ViT-B/16) Cherti et al. (2023) to extract full-resolution semantic maps. For generative tasks (e.g., scene captioning), we adopt SigLIP Zhai et al. (2023) as its visual encoder. Specifically, we extract sparse, patch-level visual embeddings from SigLIP and align them with our grid-sampled pixels using nearest-neighbor interpolation.
At runtime, we maintain a sliding window of $N_{win}=8$ keyframes, allocating $150$ optimization iterations for the radiance (geometry and appearance) phase, followed by $50$ iterations dedicated solely to the semantic phase. The entire X-GS framework operates efficiently on a single NVIDIA V100 GPU.
4.2 Contrastive VLM Results
To evaluate the visual and semantic fidelity of our system, we present qualitative comparisons of scene reconstruction and open-vocabulary object localization in Figure 3. First, we observe that the proposed X-GS-Perceiver module maintains exceptional geometry and appearance tracking; the rendered RGB reconstructions exhibit high visual fidelity and closely match the ground truth (GT) images. Second, we evaluate the distilled semantic field. Despite running in an online fashion at real-time speeds, our framework successfully distills semantic information from the VFMs of SAM and CLIP. The integration of the VQ module and grid-sampled supervision effectively balances speed and quality. Finally, we demonstrate the practical utility of our semantically enriched 3D scenes. By querying the distilled semantic field using text prompts, our framework computes per-Gaussian similarity scores to accurately localize specific objects in 3D space. This confirms that X-GS operates not merely as an efficient SLAM system, but successfully constructs a deeply comprehensible, queryable 3D semantic field ready for complex downstream tasks.
4.3 Generative VLM Results
To further demonstrate the multimodal generation capabilities of our framework, Figure 4 illustrates qualitative results for 3D scene captioning. By instantiating LLaVA-OneVision Li et al. (2025a)—backed by the Qwen2-7B-Instruct LLM Yang et al. (2024)—as the X-GS-Thinker, our system directly ingests the semantically enriched 3D Gaussians reconstructed by the X-GS-Perceiver. As shown in the figure, the Thinker successfully leverages these spatial and semantic features to generate coherent, natural language descriptions of the environment, accurately capturing both individual object properties and complex global scene layouts.
Figure 4: Qualitative results of X-GS for 3D scene caption generation.
4.4 Computational Footprint Analysis
To evaluate the efficiency and resource requirements of our system, we present a computational footprint analysis of the X-GS-Perceiver in Table 2. The table breaks down the processing duration across key pipeline components during keyframe optimization, including the time allocated for geometry, appearance, and semantic updates. Additionally, we report the overall system throughput and GPU memory load. Thanks to our highly parallelized scheduling, the heavy optimization tasks run efficiently in the background, allowing the framework to maintain real-time tracking performance on a single consumer-grade GPU.
Table 2: Computational footprint analysis of X-GS-Perceiver. We employ SigLIP and evaluate the pipeline on a single NVIDIA V100 GPU.
Main Component
Geometry & Appearance Optimization 14.0s
Semantic Supervision Target Preparation
   – Vision Encoding 0.2s
   – VQ Codebook Update 2.3s
   – Grid-Sampled Target Prefetching 4.3s
Semantic Optimization 11.2s
Average Time per Keyframe 25.2s
Average Time per Frame 2.8s
Average FPS 21.4
GPU Memory Load $\sim$9 GB
5 Conclusion
In this paper, we introduced X-GS, an extensible framework that unifies isolated 3DGS domains—including pose-free tracking, online SLAM, and semantic 3DGS—into a single cohesive system. To achieve real-time performance, we proposed the X-GS-Perceiver, which overcomes computational bottlenecks using an online VQ module with EMA updates, GPU-accelerated grid sampling, and highly parallelized scheduling. Furthermore, we demonstrated the framework’s capability for multimodal reasoning through the X-GS-Thinker, a module that bridges our semantic 3D representations with downstream VLMs and VLAs. By natively supporting complex tasks such as open-vocabulary 3D object detection, scene captioning, and embodied AI planning, X-GS establishes a robust, modular foundation for real-time spatial computing.
Limitations and Future Work
While X-GS establishes a highly extensible framework, its multimodal reasoning capabilities require more comprehensive evaluation across diverse embodied and 3D vision-language benchmarks. Currently, the interface with the X-GS-Thinker operates in a modular fashion; future work will explore end-to-end fine-tuning of the downstream VLMs natively within the 3DGS representations. Additionally, as demonstrated by recent systems like FreeSplat Wang et al. (2024b) and SplatTalk Thai et al. (2025), there is a growing trend toward feed-forward 3DGS generation and reasoning. Integrating such feed-forward mechanisms into X-GS could significantly reduce the current optimization overhead. Furthermore, leveraging emerging 3D vision foundation models—such as DUSt3R Wang et al. (2024a), MASt3R Leroy et al. (2024), or VGGT Wang et al. (2025)—could provide stronger geometric priors for mapping, further enhancing the robustness of our pose-free tracking pipeline. Finally, accommodating dynamic scenes using techniques such as 4DGS represents a highly promising direction for real-world embodied AI applications.
Appendix A 3D Gaussian Splatting
3D Gaussian Splatting (3DGS) represents a 3D scene using a collection of anisotropic 3D Gaussians, operating as a point-based alternative to continuous neural implicit representations. Each Gaussian $i$ in the comprehensive map parameter set $\Theta$ is explicitly defined by the following attributes:
  • Mean vector (Position): $\mu_{i}\in\mathbb{R}^{3}$
  • Covariance matrix: $\Sigma_{i}$
  • Opacity: $\alpha_{i}\in[0,1]$
  • Color features: $c_{i}$ (normally represented as Spherical Harmonics for view-dependence, though often simplified to base RGB values in dense SLAM configurations).
The unnormalized spatial distribution of a 3D Gaussian evaluated at a spatial point $x$ is defined as:
$G(x)=\exp\left(-\frac{1}{2}(x-\mu_{i})^{T}\Sigma_{i}^{-1}(x-\mu_{i})\right)$ (19)
To maintain positive semi-definiteness during gradient-based optimization, the covariance matrix $\Sigma_{i}$ is parameterized by a scaling matrix $S_{i}$ and a rotation matrix $R_{i}$:
$\Sigma_{i}=R_{i}S_{i}S_{i}^{T}R_{i}^{T}$ (20)
In order to render these primitives onto the image plane from a given camera pose, the 3D Gaussians are "splatted" to 2D. Given a learned viewing transformation $W$ and the Jacobian $J$ of the affine approximation of the projective alignment, the resulting 2D covariance matrix $\Sigma^{\prime}_{i}$ is calculated as:
$\Sigma^{\prime}_{i}=JW\Sigma_{i}W^{T}J^{T}$ (21)
This continuous state update sets a framework perfectly suited for dense tracking and mapping.
Appendix B 3DGS-SLAM Objectives
Continuous optical depth models compute the expected color $C$ along a camera ray by integrating transmittance $T(t)$, volume density $\sigma(t)$, and color $c(t)$. 3DGS discretizes this continuous integral using a point-based approximation. For a set of $N$ sorted 3D Gaussians projected onto the 2D image plane given a camera pose $\mathbf{T}\in\text{SE}(3)$, the color at pixel $x$ evaluates to:
$C(x\mid\mathbf{T},\Theta)=\sum_{i=1}^{N}c_{i}\alpha^{\prime}_{i}\prod_{j=1}^{i-1}(1-\alpha^{\prime}_{j})$ (22)
where $\alpha^{\prime}_{i}$ is the 2D footprint opacity. Depth $D(x)$ is similarly rendered by substituting color $c_{i}$ with the ray depth $d_{i}$.
Camera tracking is performed via Analysis-by-Synthesis. For an incoming frame at timestamp $t$, the pose is initialized via a constant velocity model. Because the 3DGS rasterizer is natively differentiable, precise tracking is achieved by freezing the map parameters $\Theta$ and running gradient descent on the camera extrinsic parameters to minimize the geometric and photometric depth losses (if depth inputs are available):
$\mathcal{L}_{\text{track}}=\mathcal{L}_{\text{photo}}(I_{t},\hat{I}_{t})+\lambda_{d}\|D_{t}-\hat{D}_{t}\|_{1}$ (23)
The global mapping thread optimizes the scene parameters $\Theta$ over a sliding window of keyframes $\mathcal{K}$ using the same combined loss, augmented with an isotropic regularization term $\mathcal{L}_{\text{iso}}$ to prevent heavily elongated artifacts:
$\mathcal{L}_{\text{map}}$ $=\sum_{k\in\mathcal{K}}\Big(\mathcal{L}_{\text{photo}}(I_{k},\hat{I}_{k})+\lambda_{d}\|D_{k}-\hat{D}_{k}\|_{1}\Big)$ (24)
$\quad+\lambda_{\text{iso}}\mathcal{L}_{\text{iso}}$
Appendix C Additional Details about Online EMA Codebook Updates
To stabilize the online EMA updates introduced in the main text, we implement active dead-code management and a warm-start initialization process.
Instead of starting from uniform counts, we initialize the EMA statistics using a confidence-filtered assignment pass over an initial buffer of semantic features. We transform the observed feature distances into confidence scores using a temperature parameter $\tau$:
$p_{nk}=\frac{\exp(-\|\mathbf{x}_{n}-\mathbf{e}_{k}\|_{2}/\tau)}{\sum_{j=1}^{K}\exp(-\|\mathbf{x}_{n}-\mathbf{e}_{j}\|_{2}/\tau)}$ (25)
Only samples where the maximum confidence exceeds a threshold $\gamma$ are allowed to contribute to the initialization of the counts $N_{k}$ and sums $\mathbf{M}_{k}$:
$m_{n}=\mathbb{I}\!\left[\max_{k}p_{nk}\geq\gamma\right]$ (26)
During the continuous online tracking stage, the confidence score masks unreliable incoming samples, while the actual codeword assignments remain hard nearest-neighbor to ensure discrete latent separation.
Furthermore, to handle dead codes escaping the tracked field of view, we explicitly monitor the EMA accumulated mass. If a codeword’s utilization falls below a dead-code threshold ($N_{k}<\delta$), it is immediately reinitialized:
$\mathbf{M}_{k}\leftarrow\tilde{\mathbf{x}},\qquad N_{k}\leftarrow 1+\varepsilon$ (27)
where $\tilde{\mathbf{x}}$ is sampled directly from a small reservoir of recently observed historical semantic features.
Appendix D Extraction of Grid-Sampled Semantic Targets
As formulated in the main text, the grid-sampled supervision relies on a target map $\mathbf{G}^{\star}$ and validity mask $\mathbf{M}$. We populate these tensors through two distinct pathways depending on the downstream semantic source:
1. Precomputed Region-Indexed Annotations: For explicit segmentation masks $S\in\{-1,0,\dots,R-1\}^{H\times W}$ (where 1-1 denotes background) and a corresponding region-feature table $\Phi\in\mathbb{R}^{R\times D}$, we directly query the table at the sampled grid location $(u,v)$:
$r_{m,n}$ $=S(u,v),$ (28)
$M_{m,n}$ $=\mathbb{I}[r_{m,n}\neq-1],$
[…]
2. Online Continuous Feature Maps: For dense semantic features streaming from an online VLM at abstract lower resolutions $\mathbf{P}\in\mathbb{R}^{D\times H_{f}\times W_{f}}$, we dynamically align the image coordinates to the low-resolution grid coordinates:
$\tilde{u}_{m,n}$ $=\operatorname{round}\!\left(u\cdot\frac{H_{f}-1}{\max(H-1,1)}\right),$ (29)
$\tilde{v}_{m,n}$ $=\operatorname{round}\!\left(v\cdot\frac{W_{f}-1}{\max(W-1,1)}\right),$
yielding the target $\mathbf{G}^{\star}_{:,m,n}=\mathbf{P}_{:,\tilde{u}_{m,n},\tilde{v}_{m,n}}$ with an unconstrained mask $M_{m,n}=1$.
Finally, to prevent the optimized point-cloud from overfitting to a static grid lattice, the offset $\mathbf{o}=(o_{h},o_{w})$ is dynamically randomized over optimization steps. By drawing the offset from the valid domain of phase shifts $\mathbf{o}_{t}\in\{0,\dots,s-1\}^{2}$, the sampled grid smoothly traverses and covers all $s^{2}$ possible pixel phases over the course of sequential mapping.
Appendix E Additional Details on Object Detection
Given a text prompt $p$, we encode it with OpenCLIP and normalize the resulting embedding:
$\mathbf{t}^{+}=\frac{E_{\text{text}}(p)}{\left\|E_{\text{text}}(p)\right\|_{2}}.$ (30)
We also encode a small set of generic negative phrases
$\mathcal{N}=\{\text{object},\ \text{things},\ \text{stuff},\ \text{texture}\},$ (31)
with normalized embeddings
$\mathbf{t}_{n}^{-}=\frac{E_{\text{text}}(n)}{\left\|E_{\text{text}}(n)\right\|_{2}},\qquad n\in\mathcal{N}.$ (32)
For semantic level $\ell$, the renderer reconstructs the feature at pixel $(u,v)$ from the rendered code weights and the learned codebook:
$\mathbf{f}_{uv}^{(\ell)}=\sum_{j=1}^{C_{\ell}}w_{uv,j}^{(\ell)}\,\mathbf{c}_{j}^{(\ell)},\qquad\hat{\mathbf{f}}_{uv}^{(\ell)}=\frac{\mathbf{f}_{uv}^{(\ell)}}{\left\|\mathbf{f}_{uv}^{(\ell)}\right\|_{2}+\varepsilon}.$ (33)
The prompt relevance at that pixel is computed by contrasting the positive query against each negative phrase:
$R_{uv}^{(\ell)}(p)=\min_{n\in\mathcal{N}}\operatorname{softmax}\!\left(\tau\begin{bmatrix}\hat{\mathbf{f}}_{uv}^{(\ell)\top}\mathbf{t}^{+}\\ \hat{\mathbf{f}}_{uv}^{(\ell)\top}\mathbf{t}_{n}^{-}\end{bmatrix}\right)_{1},$ (34)
where $(\cdot)_{1}$ denotes the probability assigned to the positive prompt and $\tau=10$ in the implementation. The final per-pixel relevance is obtained by taking the strongest response across levels:
$R_{uv}(p)=\max_{\ell}R_{uv}^{(\ell)}(p).$ (35)
In the prompt-rendering variant, the same idea is also applied directly to individual Gaussians. If Gaussian $i$ has decoded semantic feature
$\mathbf{g}_{i}=\sum_{j=1}^{C}a_{ij}\,\mathbf{c}_{j},\qquad\hat{\mathbf{g}}_{i}=\frac{\mathbf{g}_{i}}{\left\|\mathbf{g}_{i}\right\|_{2}+\varepsilon},$ (36)
its prompt relevance score is
$r_{i}(p)=\min_{n\in\mathcal{N}}\operatorname{softmax}\!\left(\tau\begin{bmatrix}\hat{\mathbf{g}}_{i}^{\top}\mathbf{t}^{+}\\ \hat{\mathbf{g}}_{i}^{\top}\mathbf{t}_{n}^{-}\end{bmatrix}\right)_{1}.$ (37)
A prompt-specific Gaussian mask is then obtained by thresholding
$m_{i}(p)=\mathbb{I}\!\left[r_{i}(p)>\delta\right].$ (38)
If no Gaussian survives the threshold, the implementation falls back to keeping the top-scoring subset.
Appendix F Entropy-Adaptive Gaussian Sampling for Generative VLMs
As demonstrated by recent works such as SplatTalk Thai et al. (2025), projecting 3D Gaussian features into discrete tokens enables Large Language Models (LLMs) to perform zero-shot 3D visual question answering and scene understanding. However, directly passing all $N_{g}$ Gaussians from an uncompressed SLAM map into the vision tower of a Generative VLM is computationally prohibitive and yields highly redundant tokens, particularly in homogeneous regions like walls and floors.
To bridge our dense semantic 3DGS map with the downstream Generative VLM, we condense the scene into a compact sequence of $M$ informative tokens ($M\ll N_{g}$) using an Entropy-Adaptive Gaussian Sampling strategy. This approach leverages the continuous-to-discrete probability distribution generated by the online Vector Quantization (VQ) module in our X-GS-Perceiver.
Let $\mathbf{f}_{i}\in\mathbb{R}^{D}$ represent the semantic feature logits of the $i$-th Gaussian over the $D$ discrete semantic clusters. We first convert these features into a semantic assignment probability distribution $p_{i,k}$ using a standard softmax operator across the feature dimension:
$p_{i,k}=\frac{\exp(f_{i,k})}{\sum_{j=1}^{D}\exp(f_{i,j})}$ (39)
The semantic entropy $H_{i}$ of the $i$-th Gaussian is then calculated as the Shannon entropy of this assignment distribution:
$H_{i}=-\sum_{k=1}^{D}p_{i,k}\log p_{i,k}$ (40)
Gaussians with a low entropy $H_{i}$ confidently match a single semantic cluster, typically representing redundant, homogeneous background surfaces. Conversely, Gaussians with a high entropy $H_{i}$ denote semantic ambiguity—indicating object boundaries, transitions between distinct instances, or complex structural information holding higher contextual value for the VLM.
To select the final $M$ tokens, we deploy a deterministic Top-$M$ sorting mechanism over the computed entropy scores. We directly extract the subset of Gaussians $\mathcal{S}$ that exhibit the highest semantic ambiguity:
$\mathcal{S}=\mathop{\operatorname{arg\,top-}M}_{i\in\{1,\dots,N_{g}\}}H_{i}$ (41)
The resulting $M$ Gaussians are then projected into the VLM’s embedding space. This strategy guarantees that the generative VLM receives an informative and tightly packed token sequence, actively filtering out redundant backgrounds while preserving critical object boundaries and structural frontiers required for robust multimodal reasoning.